OpenStudy (anonymous):

\[x ^{-2}-4x ^{-1}=10\]

OpenStudy (anonymous):

that is the same as 1/x^2-4/x=10 multiply with x^2 to get rid of the fraction 1-4x=10x^2 10x^2+4x-1=0 that has ugly solutions... \[(-4\pm \sqrt{56})/20\]

OpenStudy (anonymous):

Using my solver I get: \[x=\pm (\sqrt(14)+2)/10\]

OpenStudy (anonymous):

that is the same

OpenStudy (anonymous):

I know. I just was trying to show that there are sometimes more than one way to write a solution. I was not trying to step on any toes.